# Multi-Sources Information Fusion Learning for Multi-Points NLOS Localization

### v2

@@ Abstract
Accurate localization of mobile terminals is a pivotal aspect of integrated sensing and communication systems. Traditional fingerprint localization methods, which infer coordinates from channel information within pre-defined rectangular areas, often face challenges due to the heterogeneous distribution of fingerprints inherent in non-line-of-sight (NLOS) scenarios. To overcome this limitation, we have developed a novel multi-source information fusion learning framework referred to as the Autosync Multi-Domain NLOS Localization (AMDNLoc). Specifically, AMDNLoc employs a two-stage matched filter fused with a target tracking algorithm and iterative centroid-based clustering to automatically and irregularly segment NLOS regions, ensuring uniform fingerprint distribution within channel state information across frequency, power, and time-delay domains. Additionally, the framework utilizes a segment-specific linear classifier array, coupled with deep residual network-based feature extraction and fusion, to establish the correlation function between fingerprint features and coordinates within these regions. Simulation results demonstrate that AMDNLoc achieves an impressive accuracy of 1.46 meters on typical wireless artificial intelligence research datasets and offers interpretability, adaptability, and scalability in various scenarios.
@@ I Introduction
As a key usage scenario for sixth-generation communications, integrated sensing and communication requires a low-latency and high-precision mobile terminal (MT) localization, especially in fields like the smart city [1], internet of vehicles [2], telemedicine [3] and so on. In dense urban environments, the localization accuracy of traditional global navigation satellite systems (GNSS) and ranging-based localization techniques is significantly limited in the multi-point non-line-of-sight (NLOS) scenarios [4, 5]. Fortunately, the advent of massive multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) systems presents a new horizon for the accurate NLOS localization since these systems uniquely characterize each point within a scene by distinct multi-path features, forwarding the fingerprint-based localization methods.
Taking channel state information (CSI) associated with multi-path features as fingerprints, fingerprint-based localization method is to estimate locations by matching real-time measurements with pre-stored fingerprints, offering advantages in terms of lower energy consumption and implementation cost [6]. However, this process is particularly challenging in outdoor multi-point NLOS scenarios. The complexity arises from the influence of various urban elements, such as buildings and scattered objects, which induce substantial heterogeneity in the fingerprint distribution. This heterogeneity leads to an uneven regional correlation between fingerprints and their corresponding locations, thereby making it difficult deducing locations from fingerprints [7].
@@
To tackle the issue, conventional approaches often involve partitioning the coverage area into multiple cells, typically arranged in a grid pattern. Each cell in this grid map is associated with a reference fingerprint for localization purposes. Presently, a considerable number of research focuses on regular grid maps [8, 9, 10, 11, 12, 13, 14, 15, 16]. This approach entails segmenting a large area into medium-sized square cells based on geographical coordinates. All the fingerprints in each cell is labeled uniformly, and the cell size is iteratively reduced until the closest fingerprint in the dataset is matched. While it’s easy to manually divide regular areas, localization accuracy is highly sensitive to the specifics of the regional divisions, such as cell size and alignment. To mitigate the problem, researchers have proposed irregular grid maps, which feature variably distributed cells. For example, [17] adapts the grid to align with the urban road network, anticipating that outdoor mobile users will predominantly follow these paths. [18] introduces clustered merging, grouping geographically close fingerprints that are statistically significant for large-scale Wi-Fi positioning systems. Additionally, [19] employs the farthest-first traversal algorithm and low-discrepancy sequences to generate irregular grid maps of received signal strength (RSS). Irregular grid maps can potentially enhance performance and location estimation precision. However, the methods for generating these maps are computationally intensive. Moreover, due to the heterogeneity issue, despite the tendency for fingerprints at adjacent positions to exhibit high similarity, existing works do not guarantee that all fingerprints within a cell, whether regular or irregular, will share this level of similarity.
Building upon the foundation of manually divided regular grid maps, various artificial intelligent (AI) techniques have been developed to establish an inverse correlation function between fingerprints and locations, capitalizing on their exceptional non-linear fitting capabilities [11, 12, 13, 14, 15, 16, 20]. For instance, [13] utilizes a deep residual network (ResNet) for coarse localization, complemented by multi-layer perceptron-based transfer learning for fine tuning. [12] introduces a CALP deep convolutional neural network (DCNN), which hierarchically refines the search area to pinpoint the closest fingerprint. Furthermore, [16] presents a deep neural network architecture incorporating the Transformer model, relying on self-attention mechanisms to process sequences of fingerprint patches. While AI techniques have significantly enhanced localization accuracy, they also introduce complexities. The need for continuous mapping of real-time measurements to smaller square grid cells entails updating numerous parameters. Additionally, if the initial coarse positioning of a fingerprint is inaccurately predicted, it can lead to cumulative errors in the final results. Moreover, despite the unique multi-path features of each location, existing approaches may not fully capture this distinctiveness, since they often rely on a single type of fingerprint—such as RSS, channel impulse response (CIR), channel frequency response (CFR), the angle-delay channel amplitude matrix (ADCAM)—among others. Experimental evidence indicates that in various datasets, a proportion of remote points share strikingly similar fingerprints within a scene, challenging the foundational assumption of independent and identically distributed (i.i.d) data. This similarity can significantly undermine the accuracy of the predicted results [21, 22, 23]. Such findings suggest the need for further refinement in these AI-driven localization methods to account for the complexity and diversity of real-world environments.
To address above challenges in large-scale outdoor multi-point NLOS localization, we propose a novel multi-source information fusion learning framework named the autosync multi-domain NLOS localization (AMDNLoc). This framework, as depicted in Fig. 1, is innovative in its automatic and irregular generation of classification regions, utilizing a fusion of frequency, power, and time-delay information from CSI for the first time. The key contributions are as follows:
  • We have developed a two-stage matched filter that identifies and extracts all parallel feature intervals related to CFR distribution, named parallel feature CFR (PFCFR). This filter classifies all CFR images according to the closest PFCFR using a target tracking algorithm.
  • We employ iterative centroid-based clustering to partition the ADCAM, considering power, angle, and delay characteristics. This process combines the classification areas from both the PFCFR and ADCAM, ensuring uniform fingerprint distribution characteristics.
  • We remove the special samples through data cleaning, and introduce a segment-specific linear classifier array, which, coupled with a deep residual network for feature extraction and fusion, inversely maps fingerprints to coordinates in respective regions while maintaining global features of the complete scene.
  • The AMDNLoc framework has achieved state-of-the-art results in both ray-tracing wireless AI research dataset (WAIR-D) [24] and DeepMIMO dataset [25]. Additionally, we have validated the visualization of classification areas and the comparison of multiple processing approaches, confirming the enhanced adaptability and scalability of AMDNLoc. The full code of this paper is provided on [26].
The rest of the paper is organized as follows: Section II discusses the channel model and formulates the heterogeneity problem. Section III delves into the AMDNLoc framework, followed by numerical results in Section IV. We conclude our findings in Section V.
Lastly, specific notations are used throughout this paper for clarity. Scalars, vectors, and matrices are represented by $a$, $\mathbf{a}$, and $\mathbf{A}$, respectively. Operations on matrices such as transpose ($\mathbf{A}^{T}$), Hermitian (conjugate transpose, $\mathbf{A}^{H}$), inverse ($\mathbf{A}^{-1}$), pseudo-inverse ($\mathbf{A^{\dagger}}$), and Frobenius norm ($\|\mathbf{A}\|_{F}$) are denoted as shown. The $(m,n)$-th entry of a matrix is indicated by $a{mn}$, and ||$\cdot$|| ⋅ | denotes the modulus. The notation $diag(\mathbf{a})$ and $Tr(\mathbf{a})$ signify the diagonal matrix and trace of $\mathbf{a}$, respectively, with $o$ representing infinitesimal of higher order.
II System Model
@@
In this section, we begin by presenting the channel model, and subsequently formulate the heterogeneity problem from the NLOS fingerprint localization perspective.
@@
The considered MIMO-OFDM system is illustrated in Fig. 2. The BS is equipped with a uniform linear array (ULA) comprising $N_{t}$ antennas. We have $m\in\{1,...,M\}$ MTs, each equipped with a single omni-directional antenna. The angle of arrival (AOA) and the physical distance between the transmit antenna and the first receive antenna associated with the $p$-th path are denoted by $\phi_{p,m}\in(0,\pi)$ and $d_{p,m}$, respectively.
The CIR vector of the $m$-th user is given by:
@@
where $j=\sqrt{-1}$, $d$ is the antenna spacing (typically $\lambda/2$ in recent MIMO systems); $\tau_{p,m}=n_{p,m}T_{s}$ represents the distinguishable propagation delay associated with the $p$-th path; $n_{p,m}$ refers to the sampled delay for the $p$-th path, and $T_{s}$ denotes the sample interval. The CFR is expressed as a sum of time-domain CIRs with varying delays:
@@
${{{\bf H}}_{m}}=\left[{{{{\bf h}}_{m,0}},{{{\bf h}}_{m,1}},\ldots,{{{\bf h}}_{m,{N_{c}}-1}}}\right].$ (4)
Figure 2: The considered MIMO-OFDM system.
Then, authors in [10] established a mapping from the CFR matrix to a sparse structure by employing DFT operations, referred to as the angle-delay channel response matrix.
@@
As we know, fingerprint is determined by the locations and a specific configuration of the propagation environment:
@@
where Fingerprint represents the channel characteristics used for localization. $\boldsymbol{p}:=[x,y]$ denotes the 2D positional coordinates of the users, and $pv$ encompasses the system parameters which remain constant during training and testing, such as frequency, antenna response, and so on.
Traditional fingerprint localization aims to inversely map the Fingerprint back to $\boldsymbol{p}$, which is the inverse process of fingerprint extraction:
@@
However, the impact of different buildings on various areas differs, so that changes in the fingerprint in different areas play different roles in affecting user locations. This leads to the issue of fingerprint heterogeneity in distribution. Therefore, the inverse process of fingerprint extraction should be expressed as:
[…]
where $cv$ represents a covariate related to the region, weather, and other factors. Our objective in this research is to eliminate the regional covariate and establish a one-to-one relationship solely between $\boldsymbol{p}$ and Fingerprint.
@@ III AMDNLoc Multi-sources Framework
In this section, we delve into the heterogeneity matched filter, the network structure design, and the detailed network training methodology within the AMDNLoc framework.
@@ III-A1 PFCFR
Match Between Categories: The objective of this stage is to macroscopically merge similar categories. We utilize Eq. 9 to evaluate the similarity between the templates derived from the first-stage classification. When the similarity between any two templates exceeds $\tau_{out}$, we amalgamate all CFR images associated with these templates into a single category. This merging process continues until there is no further change in the number of final classifications. The comprehensive algorithm for this stage is methodically outlined in Alg. 2.
@@ III-A2 ADCAM
To assist in the pre-classification of ADCAM, we employ the iterative centroid-based clustering method to partition information such as the AOD, AOA, gains, and pathloss of each path into K clusters based on their distances from K class centers. One major challenge is determining the appropriate value of K. Often, it is chosen based on industry experience and may not necessarily reflect the true number of clusters present in the data. To address this challenge, we combine the Silhouette coefficients (SC) [27] and Calinski-Harabasz (CH) methods [28] to combine both the direct distances and covariance. This helps avoid scenarios where the graph of the clustering quality measure exhibits a smooth pattern (e.g., horizontal or continuously ascending/descending), making it difficult to ascertain the ideal values [29].
SC method: We calculate the average silhouette for each candidate value. The value that corresponds to the highest average silhouette score, denoted as $\mathbf{S_{SC}(i)}$, is considered the optimal number of clusters for the unsupervised learning algorithm. The silhouette score for a single data point, $S_{SC}(i)$, is computed as:
${\mathbf{S_{SC}(i)}}=\frac{b(i)-a(i)}{max\{a(i),b(i)\}},$ (11)
where ${{a(i)}}=\frac{1}{n-1}\sum_{j\neq i}^{n}distance(k(i),k(j))$ gives a measure of how well assigned the $i$-th data point is to it’s cluster, and ${{b(i)}}=\mathop{\min}_{j\neq i}\frac{1}{n}\sum^{n}distance(k(i),k(j))$ gives the average dissimilarity to the closest cluster which is not it’s cluster, ${{k(i)}}=[AOD_{i},AOA_{i},gains_{i},pathloss_{i}]$.
CH method: The CH index serves as a measure of how similar an object is to its own cluster (cohesion) relative to other clusters (separation). Cohesion is estimated based on the distances between data points within a cluster and their cluster centroid, while the separation is computed based on the distances between cluster centroids and the global centroid.
Unlike the SC method, which focuses on direct distances, the CH method considers the covariance for clustering quality. The CH index, denoted as $\mathbf{S_{CH}(k)}$, is calculated as follows:
${\mathbf{S_{CH}(k)}}=\frac{tr(B_{k})(q-k)}{tr(W_{k})(k-1)},$ (12)
where $B_{k}$/$W_{k}$ is the covariance matrix between/within categories with $k$ categories, and $q$ is the number of samples in the training set.
@@ III-A3 Data Fusing and Cleansing
To ensure high-quality input data for our model, we combine the classification results obtained from both the PFCFR and ADCAM. This integration step helps us create a comprehensive and unified classification scheme. Here’s how it works:
Combining Classifications: our initial step involves utilizing the two classification regions to refine the final classification categories. Subsequently, we apply data cleansing techniques to remove any anomalous fingerprints that might skew the analysis. In the phase of combining classifications, we adopt a systematic approach. For instance, if there are only Class 0 and Class 5 ADCAM images found within the Class 0 CFR category, we assign values of [0, 0] and [0, 5] to the new categories as Class 0 and 1, respectively. This reclassification process is methodically applied across all categories.
Data Cleansing: As previously mentioned, there may be special fingerprints in the scene that pose challenges that traditional methods struggle to handle. After our pre-processing, we can apply data cleansing techniques to address this issue. We set a number to eliminate categories with a sample size of less than or equal to the number. The rationale behind this decision is that such a small sample size in a category typically indicates an aberration, possibly due to multi-path features that significantly diverges from neighboring points. This could be attributed to erroneous recordings or the point being situated in a distinctly unique location.
@@ III-B Network Training
where ${\mathbf{\mathcal{G}_{1}}}$ and ${\mathbf{\mathcal{G}_{2}}}$ represent the mappings built by ResNet18 [30] from CFR or ADCAM images to a vector with a flattened length of $M_{H}$ or $M_{A}$ respectively. $\mathbf{\mathcal{P}}$ is a function that merges these two sets of features along the same dimensions. Finally, $\mathbf{\mathcal{G}_{3,i}}$ is a parallel multi-head mechanism in which the combined features are divided into different linear classifiers for regression operations. The variables $\sum_{i}i=I$ denote the number of final classifications, and each linear classifier corresponds to a specific region denoted as $i$. $\sum_{i}\sum_{j}i\times j=M$, and ${i,j}$ represents the $j$-th sample in the $i$-th class of the fingerprint after the heterogeneity matching.
@@
Here, $n=16$ indicates the mini-batch size.
@@ IV Numerical Results
In this section, we utilized both the WAIR-D and DeepMIMO datasets to validate the effectiveness, interpretability, and scalability of the AMDNLoc framework across various scenarios. Our approach was trained for 150 epochs with the Adam optimizer with a learning rate of $3\times 10^{-3}$ in the Nvidia 3090. Weight initialization is performed using Xavier initialization, and the Adam optimizer is utilized. The initial learning rate is set to 0.003. Network architectures and hyperparameters were chosen via cross-validation for consistency across experiments unless stated otherwise.

### v3

@@ Abstract
Accurate localization of mobile terminals is crucial for integrated sensing and communication systems. Existing fingerprint localization methods, which deduce coordinates from channel information in pre-defined rectangular areas, struggle with the heterogeneous fingerprint distribution inherent in non-line-of-sight (NLOS) scenarios. To address the problem, we introduce a novel multi-source information fusion learning framework referred to as the Autosync Multi-Domain NLOS Localization (AMDNLoc). Specifically, AMDNLoc employs a two-stage matched filter fused with a target tracking algorithm and iterative centroid-based clustering to automatically and irregularly segment NLOS regions, ensuring uniform fingerprint distribution within channel state information across frequency, power, and time-delay domains. Additionally, the framework utilizes a segment-specific linear classifier array, coupled with deep residual network-based feature extraction and fusion, to establish the correlation function between fingerprint features and coordinates within these regions. Simulation results demonstrate that AMDNLoc significantly enhances localization accuracy by over 55% compared with traditional convolutional neural network on the wireless artificial intelligence research dataset.
@@ I Introduction
As a key usage scenario for sixth-generation communication, integrated sensing and communication requires a low-latency and high-precision mobile terminal (MT) localization, especially in fields like the smart city, internet of vehicles, telemedicine and so on. Among various methods, fingerprint-based localization, leveraging unique multi-path features at each location to infer specific positions, stands out in massive multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) systems.
In particular, it treats channel state information (CSI) of multi-path features as fingerprints, and matches real-time measurements with stored fingerprints to estimate locations. However, this process, especially in outdoor multi-point non-line-of-sight (NLOS) scenarios, is fraught with complexities. The multi-path features that are influenced by diverse urban elements like buildings and scattered objects lead to substantial heterogeneity in fingerprint distribution. This results in an uneven regional on the relationship between fingerprints and locations, complicating the inverse problem of location inference from fingerprints [1].
To tackle the issue, traditional solutions involve dividing a large area into grid cells, assigning labels within each cell, and iteratively refining the search until the closest fingerprint in the dataset is matched [2, 3, 4, 5, 6, 7, 8, 9, 10]. Although manual division of areas is easy to operate, the localization accuracy can be sensitive to the specifics of regional divisions. Furthermore, due to the heterogeneity issue, while fingerprints of adjacent positions are more likely to exhibit high similarity, existing works do not guarantee that all fingerprints within a cell share a high similarity, no matter the cell is regular or not.
Building upon the manually divided areas, several artificial intelligent (AI) techniques have been proposed to build the inverse correlation function between fingerprints and locations [5, 6, 7, 8, 9, 10, 11]. However, these approaches refine the search area hierarchically which means that there are amounts of parameters to be updated and once the coarse position of the fingerprint is predicted incorrectly, errors will accumulate to the final results. What’s more, though each location’s multi-path features are unique, these approaches may fail to fully encapsulate this uniqueness by only choosing one of the received signal strength (RSS), channel impulse response (CIR), channel frequency response (CFR), and the angle-delay channel amplitude matrix (ADCAM) as the fingerprint. Experiments suggest that there exists a proportion of remote points sharing very similar fingerprints within a scene in different dataset, thereby violating the foundational assumption of independent and identically distributed (i.i.d) data and undermining the predicted results [12, 13].
@@
To solve above-mentioned issues, we propose a novel multi-source information fusion learning framework for large scale outdoor multi-points NLOS localization called the autosync multi-domain NLOS localization (AMDNLoc) framework, which automatically and irregularly generates the classification regions based on the fusion of frequency, power and time-delay information of CSI for the first time. Specifically, we develop a two-stage matched filter to identify and extract all parallel feature intervals related to CFR distribution, termed parallel feature CFR (PFCFR), and classify all CFR images according to the closest PFCFR by target tracking algorithm. Then, we utilize the iterative centroid-based clustering to partition the ADCAM considering power, angle, and delay characteristics, and combine these two classification areas maintaining the same fingerprint distribution characteristics. Then, special samples are eliminated by data cleaning and a segment-specific linear classifier array coupled with deep residual network-based feature extraction and fusion is introduced to inversely map the fingerprints to the coordinates in the respective regions while maintaining the global features. State of the art (SOTA) results are achieved in the ray-tracing wireless AI research dataset (WAIR-D) [14, 15]. Also the visualization of classification areas and comparison of multiple processing are validated to ensure the enhanced adaptability and scalability of AMDNLoc. The code is avalable in the [16].
II System and Channel Model
@@
In this section, we begin by presenting the channel model and subsequently formulate the heterogeneity problem in NLOS fingerprint localization.
@@
The channel transmission diagram in a MIMO-OFDM system is illustrated in Fig. 2. The BS is equipped with a uniform linear array (ULA) comprising $N_{t}$ antennas. We have $m\in\{1,...,M\}$ MTs, each equipped with a single omni-directional antenna. The angle of arrival (AOA) and the physical distance between the transmit antenna and the first receive antenna associated with the $p$-th path are denoted by $\phi_{p,m}\in(0,\pi)$ and $d_{p,m}$, respectively.
The CIR vector associated with the $p$-th path of the $m$-th user is given by:
@@
where $j=\sqrt{-1}$, $d$ is the antenna spacing (typically $\lambda/2$ in recent MIMO systems); $\tau_{p,m}=n_{p,m}T_{s}$ represents the distinguishable propagation delay associated with the $p$-th path, $n_{p,m}$ refers to the sampled delay for the $p$-th path, and $T_{s}$ denotes the sample interval. The CFR is expressed as a sum of time-domain CIRs with varying delays:
@@
${{{\bf H}}_{m}}=\left[{{{{\bf h}}_{m,0}},{{{\bf h}}_{m,1}},\ldots,{{{\bf h}}_{m,{N_{c}}-1}}}\right],$ (4)
Figure 2: The wireless channel from MT $m$ to the BS.
Then, authors in [4] established a mapping from the CFR matrix to a sparse structure by employing DFT operations, referred to as the angle-delay channel response matrix. It is a linear transformation of the CSI computed by multiplying it with two DFT matrices.
@@
As we know, fingerprint is determined by the locations and a specific configuration of the propagation environment, given by:
@@
where Fingerprint represents the channel characteristics used for localization. $\boldsymbol{p}:=[x,y]$ denotes the 2D positional coordinates of the users, and $pv$ encompasses the system parameters which remain constant during training and testing.
Traditional fingerprint localization aims to inversely map the Fingerprint back to $\boldsymbol{p}$. However, the impact of different buildings on various areas differs, so that changes in the fingerprint in different areas play different roles in affecting user locations. This leads to the issue of fingerprint heterogeneity in distribution. Therefore, the inverse process of fingerprint extraction should be expressed as:
@@
where $cv$ represents a covariate related to the region, weather, and other factors. Our objective in this research is to eliminate the regional covariate from fingerprint localization and establish a one-to-one relationship solely between $\boldsymbol{p}$ and Fingerprint.
@@ III AMDNLoc Multi-sources Framework
In this section, we delve into the heterogeneity matched filter, the design of the network structure, and the detailed network training methodology within the AMDNLoc framework.
@@ III-A1 PFCFR
Match Between Categories: The objective of this stage is to macroscopically merge similar categories. We utilize Eq. 8 to evaluate the similarity between the templates derived from the first-stage classification results. When the similarity measurement between any two templates exceeds $\tau_{out}$, we proceed to amalgamate all the CFR images associated with these templates into a single category. This merging process continues until there is no further change in the number of final classifications. The comprehensive algorithm for this stage is methodically outlined in Alg. 2.
@@ III-A2 ADCAM
To assist in the pre-classification of ADCAM, we employ the iterative centroid-based clustering method to partition information such as the AOD, AOA, gains, and pathloss of each path into K clusters based on their distances from K class centers. To choose the appropriate value of K, we combine the Silhouette coefficients and Calinski-Harabasz methods to combine both the direct distances and covariance [17]. This helps avoid scenarios where the graph of the clustering quality measure exhibits a smooth pattern (e.g., horizontal or continuously ascending/descending), making it difficult to ascertain the ideal K value.
@@ III-A3 Data Fusing and Cleansing
To ensure high-quality input data for our model, our initial step involves utilizing the two classification regions to refine the final classification categories. Subsequently, we apply data cleansing techniques to remove any anomalous fingerprints that might skew the analysis. In the phase of combining classifications, we adopt a systematic approach. For instance, if there are only Class 0 and Class 5 ADCAM images found within the Class 0 CFR category, we assign values of [0, 0] and [0, 5] to the new categories as Class 0 and 1, respectively. This reclassification process is methodically applied across all categories.
During the data cleansing phase, we eliminate any category that contains a sample size of 2 or fewer. The rationale behind this decision is that such a small sample size in a category typically indicates an aberration, possibly due to multi-path features that significantly diverges from neighboring points. This could be attributed to erroneous recordings or the point being situated in a distinctly unique location.
@@ III-B Network Training
where ${\mathbf{\mathcal{G}_{1}}}$ and ${\mathbf{\mathcal{G}_{2}}}$ represent the mappings built by deep residual network 18 (ResNet18 [18]) from CFR or ADCAM images to a vector with a flattened length of $M_{H}$ or $M_{A}$ respectively. $\mathbf{\mathcal{P}}$ is a function that merges these two sets of features along the same dimensions. Finally, $\mathbf{\mathcal{G}_{3,i}}$ is a parallel multi-head mechanism in which the combined features are divided into different linear classifiers for regression operations. The variables $\sum_{i}i=I$ denote the number of final classifications, and each linear classifier corresponds to a specific region denoted as $i$. $\sum_{i}\sum_{j}i\times j=M$, and ${i,j}$ represents the $j$-th sample in the $i$-th class of the fingerprint after the heterogeneity matching.
@@
Here, $n=16$ indicates the mini-batch size. Weight initialization is performed using Xavier initialization, and the Adam optimizer is utilized. The initial learning rate is set to 0.003.
Figure 4: NLOS pre-classification.
@@ IV Numerical Results
This section demonstrates our framework’s performance through simulations on scenario 00743 of the WAIR-D dataset. Our approach was trained for 150 epochs with the Adam optimizer with a learning rate of $3\times 10^{-3}$ in the Nvidia 3090. Network architectures and hyperparameters were chosen via cross-validation for consistency across experiments unless stated otherwise.